Can anyone help me with these problems???

1) Use the rules of exponents to evaluate or simplify. Write without negative exponents.
y^-5 / y^-2 = ____

2) Use the rules of exponents to evaluate or simplify. Write without negative exponents.
(-243)^3/5 = ____

3) Use the rules of exponents to evaluate or simplify. Write without negative exponents.
36^1/2 = ____

4) Simplify by expressing fractional exponents instead of radicals.

sqrt ab

5) Simplify by expressing fractional exponents instead of radicals.

8th root of x^2y^4

6) Simplify:
8th root of y^2

7) Simplify:
5 / sqrt of 5

8) Simplify:
3rd root of 27^2

Thank You for your help. :)
~Sarah

Answer by stanbon(75887)

1) Use the rules of exponents to evaluate or simplify. Write without negative exponents.
y^-5 / y^-2 = 1/y^(-2--5) = 1/y^3
---------------------------------------------
2) Use the rules of exponents to evaluate or simplify. Write without negative exponents.
(-243)^(3/5) = (-3)^3 = -27
--------------------------------------------------
3) Use the rules of exponents to evaluate or simplify. Write without negative exponents.
36^1/2 = 6
-------------------------------------------------
4) Simplify by expressing fractional exponents instead of radicals.
sqrt ab = (ab)^(1/2)
--------------------------------------------------
5) Simplify by expressing fractional exponents instead of radicals.
8th root of x^2y^4
= (x^2y^4)^(1/8) = x^(1/4)*y^(1/2)
--------------------------------------------------
6) Simplify:
8th root of y^2
= (y^2)^(1/8)
= y^(1/4)
---------------------------------------------------
7) Simplify:
5 / sqrt of 5
What do you mean?
---------------------------------------------------
8) Simplify:
3rd root of 27^2
= (27)^(2/3)
= (3)^2
= 9
====================================================
Cheers,
Stan H.